OpenStudy (anonymous):

If (a + 3) / 5 is an integer, what is remainder when ‘a’ is divided by 5?

OpenStudy (pratyush5):

i didnt solve it but by hit and trial it is 2

OpenStudy (amistre64):

(a+3)/5 = n a+3 = 5n a = 5n - 3 a = -3 mod 5; or simply 2 mod 5